OpenStudy (anonymous):

HELP PLEASE!! Solve the equation for y. 8x – 9y = 11 A. y=-8/9x-9/11 B. y=9/8x+11/8 C. y=-8x-11 D. y=8/9x-11/9

OpenStudy (nethercreep333):

Let's solve for x. 8x−9y=11 Step 1: Add 9y to both sides. 8x−9y+9y=11+9y 8x=9y+11 Step 2: Divide both sides by 8. 8x 8 = 9y+11 8 x= 9 8 y+ 11 8

OpenStudy (nethercreep333):

mazbe b

OpenStudy (anonymous):

@Vocaloid can you help?

OpenStudy (anonymous):

nevermind @Vocaloid